Solve each part of this problem using a Proportion.
Use the variable x to represent the unknown number of cups of sugar. A recipe requires 1 1/3 cups of sugar to make 8 giant muffins. How many cups of sugar are needed for 36 giant muffins?

Answer by Theo(13342)

1 + 1/3 cups of suger for 8 giant muffins.
how many for 36?

36/8 = 4.5

4.5 * 1 + 1/3 = 4.5 * 4/3 = 6

you would need 6 cups of sugar for 36 giant muffins.

the ratio you set up is:

(4/3) / 8 = x / 36

you cross multiply to get:

(4/3) * 36 = 8 * x

you simplify to get:

144 / 3 = 8 * x

divide both sides of the equation by 8 to get:

(144/3)/8 = x

simplify to get:

144/(3*8) = x

simplify further to get:

144/24 = x

solve for x to get x = 144/24 = 6